OpenStudy (anonymous):

evalute: 3 - 2/(3-(2/3))

OpenStudy (anonymous):

15/7

OpenStudy (callisto):

\[3-\frac{2}{3-(2-\frac{2}{3})}\]\[=3-\frac{2}{3-(\frac{4}{3})}\]\[=3-\frac{2}{(\frac{5}{3})}\]\[=3-\frac{6}{5} = \frac{9}{5}\]

OpenStudy (anonymous):

thank for showing the steps callisto

OpenStudy (anonymous):

@tim538 its wrong callisto

OpenStudy (anonymous):

15/7 is the answer

OpenStudy (anonymous):

@Callisto you add 2 to the equation w/c is wrong

OpenStudy (anonymous):

so 15/7 is the right answer?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Ah... My bad :| I'm sorry!!!!! \[3-\frac{2}{3-\frac{2}{3}} = 3-\frac{2}{\frac{7}{3}}= 3-\frac{6}{7}=\frac{15}{7}\] @traile is correct!!!!!

OpenStudy (anonymous):

thank traile